# Uncertainty-Aware Surrogate-based Amortized Bayesian Inference for Computationally Expensive Models

### v2

@@ 4.4.2 Validation of Surrogate Model and Parameter Inference Setup
Since inference will be performed using real measurement data, the first step is to evaluate how accurately the surrogate represents the underlying system. Figure˜7(a) presents the measurement data, the model evaluations used for surrogate training, and the surrogate predictions obtained from $10^{4}$ prior samples across the eight locations. The results indicate that the simulated model data—and consequently the surrogate—do not cover the region occupied by the measurement data. This makes the measurement data out-of-distribution (OOD) relative to the simulation model’s a priori predictions - a phenomenon that is suspected to occur frequently in dynamic systems modeling, although rarely investigated in detail.
@@
Neural networks are known to generalize poorly outside their training domain, resulting in poor posteriors obtained via ABI (Schmitt et al., 2024a), which is shown in Section˜E.1. Thus, we seek to transfer this challenge to the surrogate instead. To this end, we found that the surrogate priors used for parameter inference can be adjusted to ensure that the measurement data fall within the surrogate’s support. This approach leverages the generalization properties of the low-dimensional, regularized surrogate, avoiding dependence on the poor generalization abilities of neural networks, which are used within the ABI model during inference.
The results of a sensitivity analysis (more detailed in Section˜E.2) on the surrogate using Sobol indices (Sobol’, 1990) shows that the surrogate’s output at all locations is almost only sensitive to $k_{ub}$.
In Fig.˜8, the surrogate output is shown exemplarily at location index 3 by varying $k_{ub}$. It can be observed that the uncertainty of the surrogate increases when evaluating OOD inputs. Based on these results, we adapt the parameter inference prior range for $k_{ub}$, such that $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-3})$. This adjustment results in a prior predictive that covers the measurement data, as illustrated in Fig.˜7(b).
@@ 4.4.3 Comparison of Parameter Posterior
Since parameter inference is carried out on real measurement data, no ground truth is available. We therefore compare only the inferred posteriors obtained using low-budget ABI ($N_{B}=25$), SABI, UA-SABI, Point, and E-Post. For easy visual comparison, we applied kernel density estimation (KDE) with Scott’s bandwidth method to the posterior samples from each approach. Since $k_{ub}$ is the only sensitive parameter, we present its posterior estimates in Fig.˜9 to avoid redundancy, while posterior estimates for the remaining parameters are provided in Section˜E.3.
@@
Figure˜9 demonstrates that low-budget ABI, which is shown as a point estimate due to quasi-nonexistent variance in the posterior, exhibits a pronounced bias and substantially underestimates posterior uncertainty. This confirms again that training an ABI model with very limited data is not meaningful. Comparing the surrogate-based methods further highlights that explicitly quantifying surrogate uncertainty and propagating it through the inference process is reflected in the wider posterior estimates. To assess the validity of the obtained posteriors and the appropriateness of the quantified and propagated uncertainty, Fig.˜10 presents the posterior predictive distributions for calcite across the four surrogate-based methods.
@@
Figure˜10(a) and Fig.˜10(b) clearly illustrate that not quantifying and propagating uncertainty results in overconfident posterior predictives. In contrast, as shown in Fig.˜10(c) and Fig.˜10(d), explicitly accounting for uncertainty produces confidence intervals that successfully cover the measurement data. Both methods yield comparable intervals, with UA-SABI producing a slightly broader interval as a higher posterior density in the out-of-distribution region of the most sensitive parameter $k_{ub}$ (Fig.˜9) yields increased surrogate predictive uncertainty.
@@ 5 Summary and Outlook
In this work, we applied UA-SABI only to low-dimensional, computationally expensive problems to keep results interpretable and avoid challenges that would distract from our scope - such challenges as they appear in surrogate training and parameter inference for high-dimensional models with only scarce available data. Future work should test the method on higher-dimensional parameter spaces to further evaluate its scalability and performance. Additionally, the surrogate is technically misspecified, as the simulator is not fully contained within the surrogate class. Thus, future work could integrate UA-SABI with recently proposed methods for detecting and mitigating model misspecification in ABI (Schmitt et al., 2024a; Dellaporta et al., 2022; Elsemüller et al., 2025), which are all naturally compatible with our framework.
@@ Acknowledgments
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding – EXC2075 – 390740016 under Germany’s Excellence Strategy and the Collaborative Research Centre SFB 1313, Project Number 327154368. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We further acknowledge the support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. Additionally, we thank Dr. Ilja Kröker for his help and and to both Dr. Kröker and Dr. Johannes Hommel for providing data.
Appendix A Pseudocode
@@
20:   Pass $(\bm{\omega}^{(i)},\mathbf{s}^{(i)})$ through inference network $q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid\mathbf{s}^{(i)})$
21:   Compute loss from Eq.˜2 and update network parameters $\bm{\theta},\bm{\varphi}$
22:  end for
@@ Appendix B Proofs
The posterior targeted by UA-SABI is the same as the E-Post posterior.
@@
In UA-SABI, as per Eq.˜8, we sample from the joint distribution
@@
to train the conditional neural density estimator $q(\bm{\omega}\mid\mathbf{y})$. We only condition on the data $\mathbf{y}$ and treat the surrogate parameters $(\mathbf{c},\sigma)$ as nuisance parameters ignored by $q(\bm{\omega}\mid\mathbf{y})$. This means that we (implicitly) integrate over the distribution of $(\mathbf{c},\sigma)$. Hence, UA-SABI targets the following posterior:
$p_{\rm UA-SABI}(\bm{\omega}\mid\mathbf{y},D_{T})\propto\iint p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma$ (15)
On the other hand, E-Post targets the following posterior, as per Eq.˜10:
$p_{\rm E-Post}(\bm{\omega}\mid\mathbf{y},D_{T})$ $=\iint p(\bm{\omega}\mid\mathbf{y},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma$ (16)
$\propto\iint p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})\,d\mathbf{c}\,d\sigma,$ (17)
@@ C.2 Neural Posterior Estimation
For the first two case studies, we used the same NPE setup. The summary network $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ is a DeepSet Zaheer et al. (2017) composed of two hidden layers, each containing 10 neurons. It outputs a 10-dimensional summary vector. For the inference network $I_{\bm{\varphi}}(\mathbf{s})$, we employ a coupling flow as implemented in Radev et al. (2023b). For the third case study, we did not use a summary network as the measurement locations in the experiment were fixed. For the inference network $I_{\bm{\varphi}}(\mathbf{s})$, we used a spline coupling flow (Durkan et al., 2019) as implemented in Radev et al. (2023b). The training process employed a cosine learning rate scheduler with an initial learning rate of $5\times 10^{-4}$ and a minimum learning rate fraction of $\alpha=10^{-6}$. The scheduler operated over a total of 12,800 steps, corresponding to 128 batches per epoch over 100 epochs. All NPE models were trained using the Adam optimizer (Kingma & Ba, 2017) for 100 epochs.
Appendix D Additional Results for Case Study 2: $\text{CO}_{2}$ Storage Model
@@
In Fig.˜11 we show that SABI and Point as well as UA-SABI and E-Post yield similar results in both recovery and calibration. The notably wide intervals observed for Point, especially compared to SABI in the recovery plots suggest convergence problems in the MCMC, likely due to the low standard deviation in the likelihood. These convergence problems may explain the inconsistencies between SABI and Point. In comparison, their uncertainty-aware counterparts UA-SABI and E-Post produce highly similar estimates.
@@
Appendix E Additional Results for Case Study 3: MICP Model
E.1 Parameter Posterior with Original Priors
Performing inference with the original prior for $k_{ub}$ leads, regardless of the inference method, to point-like posteriors at the upper boundary of the prior. The reason is the same for all methods: the maximal value of $k_{ub}$ yields the highest output values. Since the measurement data are always higher than the surrogate output under the original prior, the posterior collapses to a point-like estimate at the maximum admissible value of $k_{ub}$, as exemplified in Fig.˜8.
@@ E.1 Parameter Posterior with Original Priors
E.2 Sensitivity Analysis
We conducted a sensitivity analysis on the surrogate to identify the parameters to which the surrogate output is most sensitive. Sobol indices (Sobol’, 1990) decompose the output variance into contributions from each parameter (including their interactions), with the total Sobol index of a parameter quantifying its overall influence on the output variance. Figure˜13 presents the distributions of the total Sobol indices (Le Gratiet et al., 2017) for each parameter at each location, noting that each location has its own surrogate. Since the Bayesian PCE yields samples of the coefficients, the corresponding Sobol indices are also obtained in the form of distributions. Figure˜13 clearly shows that almost all the variance in the output at all locations is determined by $k_{ub}$.
@@ E.2 Sensitivity Analysis
E.3 Parameter Posterior for Remaining Parameters
Figure˜14 shows the KDEs of 4,000 parameter samples for the 3 remaining parameters, $c_{a,1},c_{a,2}$, and $\rho_{f}$, obtained with low-budget ABI, Point, E-Post, SABI, and UA-SABI. Two distinct scenarios emerge: 1) the posterior collapses to a point estimate (often at the boundary of the prior), or 2) the posterior remains close to the prior distribution. Point-like posteriors are produced by low-budget ABI or Point, while surrogate-based methods yield posteriors resembling the prior. Both outcomes can be explained by the lack of sensitivity of the output to these parameters, as shown by the total Sobol indices in Fig.˜13. This missing sensitivity makes reliable inference very difficult.
Figure 14: KDEs of 4,000 parameter posterior samples for the remaining parameters $c_{a,1},c_{a,2}$, and $\rho_{f}$ of the MICP model obtained with low-budget ABI, Point, E-Post, SABI and UA-SABI.
E.4 Runtime Comparison

### v3

@@ 4.4.2 Validation of Surrogate Model and Parameter Inference Setup
Since inference will be performed using real measurement data, the first step is to evaluate how accurately the surrogate represents the underlying system. Figure˜8(a) presents the measurement data, the model evaluations used for surrogate training, and the surrogate predictions obtained from $10^{4}$ prior samples across the eight locations. The results indicate that the simulated model data—and consequently the surrogate – do not cover the region occupied by the measurement data. This makes the measurement data out-of-distribution (OOD) relative to the simulation model’s a priori predictions – a phenomenon that is suspected to occur frequently in dynamic systems modeling, although rarely investigated in detail.
@@
Neural networks are known to generalize poorly outside their training domain, resulting in poor posteriors obtained via ABI (Schmitt et al., 2024a), which is shown in Section˜F.2. Thus, we seek to transfer this challenge to the surrogate instead. To this end, we found that the surrogate priors used for parameter inference can be adjusted to ensure that the measurement data fall within the surrogate’s support. This approach leverages the generalization properties of the low-dimensional, regularized surrogate, avoiding dependence on the poor generalization abilities of neural networks, which are used within the ABI model during inference.
The results of a sensitivity analysis (more detailed in Section˜F.3) on the surrogate using Sobol indices (Sobol’, 1990) shows that the surrogate’s output at all locations is almost only sensitive to $k_{ub}$.
In Fig.˜9, the surrogate output is shown exemplarily at location index 3 by varying $k_{ub}$. It can be observed that the uncertainty of the surrogate increases when evaluating OOD inputs. Based on these results, we adapt the parameter inference prior range for $k_{ub}$, such that $k_{ub}\sim\mathcal{U}(10^{-5},5\cdot 10^{-3})$. This adjustment results in a prior predictive that covers the measurement data, as illustrated in Fig.˜8(b).
@@ 4.4.3 Comparison of Parameter Posterior
Since parameter inference is carried out on real measurement data, no ground truth is available. We therefore compare only the inferred posteriors obtained using low-budget ABI ($N_{B}=25$), SABI, UA-SABI, Point, and E-Post. For easy visual comparison, we applied kernel density estimation (KDE) with Scott’s bandwidth method to the posterior samples from each approach. Since $k_{ub}$ is the only sensitive parameter, we present its posterior estimates in Fig.˜10 to avoid redundancy, while posterior estimates for the remaining parameters are provided in Section˜F.4.
@@
Figure˜10 demonstrates that low-budget ABI, which is shown as a point estimate due to quasi-nonexistent variance in the posterior, exhibits a pronounced bias and substantially underestimates posterior uncertainty. This confirms again that training an ABI model with very limited data is not meaningful. Comparing the surrogate-based methods further highlights that explicitly quantifying surrogate uncertainty and propagating it through the inference process is reflected in the wider posterior estimates. However, E-Post and UA-SABI posteriors exhibit slight difference, which we attribute to non-convergence of some MCMC runs in E-Post. To assess the validity of the obtained posteriors and the appropriateness of the quantified and propagated uncertainty, Fig.˜11 presents the posterior predictive distributions for calcite across the four surrogate-based methods.
@@
Figure˜11(a) and Fig.˜11(b) clearly illustrate that not quantifying and propagating uncertainty results in overconfident posterior predictives. In contrast, as shown in Fig.˜11(c) and Fig.˜11(d), explicitly accounting for uncertainty produces confidence intervals that successfully cover the measurement data. Both methods yield comparable intervals, with UA-SABI producing a slightly broader interval as its higher posterior density in the out-of-distribution region of the most sensitive parameter $k_{ub}$ (Fig.˜10) yields increased surrogate predictive uncertainty.
@@ 5 Summary and Outlook
In this work, we applied UA-SABI only to low-dimensional, computationally expensive problems to keep results interpretable and avoid challenges that would distract from our scope. While this focus allowed us to clearly demonstrate the method’s core advantages, the question of scalability to higher dimensions naturally arises. In general, standard ABI performs well in high-dimensional settings when sufficient simulated training data is available (e.g. Zhou et al., 2025). However, data-scarcity necessitates the use of surrogate models to augment the limited training data. Extending UA-SABI to high-dimensional problems thus hinges on developing surrogates that satisfy two critical requirements: robustness to high-dimensionality and minimal training data requirements. If these conditions cannot be met, standard ABI remains the more practical choice. High-dimensional surrogate modeling, however, presents fundamental challenges. In the PCE framework, for example, a high-dimensional parameter space results in a high-dimensional coefficient space and thus a high-dimensional integration problem, which suffers from the curse of dimensionality. Consequently, the surrogate type must be chosen to accommodate this challenge. Neural network surrogates could be one option, but they require substantial training data again (similar to standard ABI) so nothing is gained here. In PCE, existing approaches try to address high-dimensionality through input dimensionality reduction techniques (e.g. Li and Tartakovsky, 2020) or sparsity methods in PCE (e.g. Bürkner et al., 2023; Luthen et al., 2021) to reduce either the parameter space or the integration space. Combining these methods would be a valuable direction for future research.
Additionally, the surrogate is technically misspecified, as the simulator is not fully contained within the surrogate class. Thus, future work could integrate UA-SABI with recently proposed methods for detecting and mitigating model misspecification in ABI (Schmitt et al., 2024a; Dellaporta et al., 2022; elsemüller2025doesunsuperviseddomainadaptation), which are all naturally compatible with our framework.
Author Contributions
This work represents an equal contribution of Stefania Scheurer and Philipp Reiser. Following the CRediT taxonomy:
Stefania Scheurer: Conceptualization, Methodology, Software, Writing - Original Draft, Writing - Review & Editing, Philipp Reiser: Conceptualization, Methodology, Software, Writing - Original Draft, Writing - Review & Editing, Tim Brünnette: Writing - Review & Editing, Wolfgang Nowak: Writing - Review & Editing, Anneli Guthke: Writing - Review & Editing, Paul-Christian Bürkner: Conceptualization, Writing - Review & Editing.
@@ Acknowledgments
We thank the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) for supporting this work by funding – EXC2075 – 390740016 under Germany’s Excellence Strategy and the Collaborative Research Centre SFB 1313, Project Number 327154368. We acknowledge the support by the Stuttgart Center for Simulation Science (SimTech). We further acknowledge the support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526. Additionally, we thank Dr. Ilja Kröker for his help and to both Dr. Kröker and Dr. Johannes Hommel for providing data.
Appendix A Algorithmic Overview
Algorithm˜1 gives an algorithmic overview of UA-SABI training. Note that this is for online training. In case of offline training, inputs and parameters $\{(\mathbf{x}^{(i)},\bm{\omega}^{(i)})\}_{i=1}^{N_{B}}$ are fixed and not resampled in each epoch.
@@ Appendix A Algorithmic Overview
20:   Pass $\mathbf{s}^{(i)}$ through inference network $I_{\bm{\varphi}}(\mathbf{s}^{(i)})$ which implies $q_{\bm{\varphi}}(\bm{\omega}^{(i)}\mid\mathbf{s}^{(i)})$
21:  end for
22:  Compute loss from Eq.˜5 and update network parameters $\bm{\theta},\bm{\varphi}$
@@ Appendix B Proofs
We refer to ABI standard conditions as those under which ABI yields asymptotically correct posteriors, following Radev et al. (2020). Specifically, we assume (i) an infinitely large training dataset, (ii) a conditional neural density estimator $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{y})$ that is sufficiently expressive, (iii) convergence of the training procedure to the true conditional density, and (iv) inference data drawn from the same data-generating process as the training data.
Likewise, MCMC standard conditions refer to ergodicity, irreducibility, and aperiodicity of the Markov chain (e.g., Robert and Casella, 2004), ensuring convergence of the chain to the target posterior distribution as the number of samples tends to infinity.
@@
Under ABI standard conditions, and assuming an infinite number of samples used to propagate from the surrogate posterior $p(\mathbf{c},\sigma\mid D_{T})$, the posterior distribution targeted by UA-SABI converges to the E-Post posterior.
@@
In UA-SABI, as per Eq.˜11, we sample from the joint distribution
@@
to train the conditional neural density estimator $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{y})$. We only condition on the data $\mathbf{y}$ and treat the surrogate parameters $(\mathbf{c},\sigma)$ as nuisance parameters ignored by $q_{\bm{\varphi}}(\bm{\omega}\mid\mathbf{y})$. This means that we (implicitly) integrate over the distribution of $(\mathbf{c},\sigma)$. Hence, UA-SABI targets the following posterior:
$p_{\rm UA-SABI}(\bm{\omega}\mid\mathbf{y},D_{T})\propto\iint p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})\,\text{d}\mathbf{c}\,\text{d}\sigma.$ (18)
On the other hand, E-Post targets the following posterior, as per Eq.˜13:
$p_{\rm E-Post}(\bm{\omega}\mid\mathbf{y},D_{T})$ $=\iint p(\bm{\omega}\mid\mathbf{y},\mathbf{c},\sigma)\,p(\mathbf{c},\sigma\mid D_{T})\,\text{d}\mathbf{c}\,\text{d}\sigma$ (19)
$\propto\iint p(\mathbf{y}\mid\bm{\omega},\mathbf{c},\sigma)\,p(\bm{\omega})\,p(\mathbf{c},\sigma\mid D_{T})\,\text{d}\mathbf{c}\,\text{d}\sigma,$ (20)
@@
Corollary 1.
For an infinite number of samples from the surrogate posterior $p(\mathbf{c},\sigma\mid D_{T})$ and under ABI and MCMC standard conditions, the empirical distributions of UA-SABI samples and MCMC samples from E-Post converge to each other.
@@ C.2 Neural Posterior Estimation
For the first two case studies, we used the same NPE setup. The summary network $S_{\bm{\theta}}(\mathbf{x},\mathbf{y})$ is a DeepSet Zaheer et al. (2017) composed of two hidden layers, each containing 10 neurons. It outputs a 10-dimensional summary vector. For the inference network $I_{\bm{\varphi}}(\mathbf{s})$, we employ a coupling flow as implemented in Radev et al. (2023b). For the third case study, we did not use a summary network as the measurement locations in the experiment were fixed. For the inference network $I_{\bm{\varphi}}(\mathbf{s})$, we used a spline coupling flow (Durkan et al., 2019) as implemented in Radev et al. (2023b). The training process employed a cosine learning rate scheduler with an initial learning rate of $5\times 10^{-4}$ and a minimum learning rate fraction of $\alpha=10^{-6}$. The scheduler operated over a total of 12,800 steps, corresponding to 128 batches per epoch over 100 epochs. All NPE models were trained using the Adam optimizer (Kingma and Ba, 2017) for 100 epochs.
Appendix D Additional Results for Case Study 1: LogSin Model
This section shows the MCMC convergence analysis of E-Post for the LogSin model to explain the differences between UA-SABI and E-Post. We follow the same setup as described in Section˜4.2.1, but used more samples and chains to check convergence. We analyzed the convergence for each of the 200 ground truth parameters and for each surrogate posterior sample respectively. We run MCMC separately for each of the 1,000 surrogate posterior samples with 4 chains and 1,000 warm-up and 1,000 sampling iterations, respectively, resulting in 1,000 $\times$ 4 $\times$ 1,000 draws per ground truth parameter used for analysis.
For each surrogate posterior sample we can check the convergence of the MCMC run using standard diagnostics $\widehat{R}$. As convergence threshold, we follow the standard recommendation $\widehat{R}<1.01$ (Vehtari et al., 2021). Figure˜12 shows the ratio of non-converged runs ($\widehat{R}$ > 1.01) started by each surrogate posterior sample (SPS) vs. total number of runs for each of the 200 ground truth parameters. It is evident that for most ground truth parameter values more than 10% of the runs have not converged, with some having more than 40% of the runs with an $\widehat{R}>1.01$. Notably, for most ground truth parameter values, more than 10% of the runs have not converged, with some parameters even exceeding more than 40% for that ratio. This indicates that the E-Post posterior is not reliable, which results in deviating posteriors for E-Post and UA-SABI, as shown in Fig.˜5.
Figure 12: Ratio of non-converged runs ($\widehat{R}$ > 1.01) started by each surrogate posterior sample (SPS) vs. total number of runs for each of the 200 ground truth parameters for the LogSin model.
Additionally, Fig.˜13 shows the trace plots of single E-Post runs (4 chains) for two exemplary draws of the surrogate posterior with bad $\widehat{R}$ given a single ground truth parameter for the LogSin model. These traces indicate that the chains fail to mix properly between the two modes of the underlying posterior. When the modes have little overlap, MCMC struggles to recover their correct proportions.
(a) Sample 1.
(b) Sample 2.
Figure 13: Trace plots of single E-Post MCMC runs for two exemplary draws of the surrogate posterior given a single ground truth parameter for the LogSin model. The different colors indicate different chains.
Appendix E Additional Results for Case Study 2: $\text{CO}_{2}$ Storage Model
@@
In Fig.˜14 we show that SABI and Point as well as UA-SABI and E-Post yield similar results in both recovery and calibration. The notably wide intervals observed for Point, especially compared to SABI in the recovery plots suggest convergence problems in the MCMC, likely due to the low standard deviation in the likelihood. These convergence problems may explain the inconsistencies between SABI and Point. In comparison, their uncertainty-aware counterparts UA-SABI and E-Post produce highly similar estimates.
@@
Appendix F Additional Results for Case Study 3: MICP Model
F.1 Illustration
Figure 15: Schematic illustration of relevant processes during MICP. Figure adapted from Scheurer et al. (2021), licensed under $C$ BY 4.0.
F.2 Parameter Posterior with Original Priors
Performing inference with the original prior for $k_{ub}$ leads, regardless of the inference method, to point-like posteriors at the upper boundary of the prior. The reason is the same for all methods: the maximal value of $k_{ub}$ yields the highest output values. Since the measurement data are always higher than the surrogate output under the original prior, the posterior collapses to a point-like estimate at the maximum admissible value of $k_{ub}$, as exemplified in Fig.˜9.
@@ F.2 Parameter Posterior with Original Priors
F.3 Sensitivity Analysis
We conducted a sensitivity analysis on the surrogate to identify the parameters to which the surrogate output is most sensitive. Sobol indices (Sobol’, 1990) decompose the output variance into contributions from each parameter (including their interactions), with the total Sobol index of a parameter quantifying its overall influence on the output variance. Figure˜17 presents the distributions of the total Sobol indices (Le Gratiet et al., 2017) for each parameter at each location, noting that each location has its own surrogate. Since the Bayesian PCE yields samples of the coefficients, the corresponding Sobol indices are also obtained in the form of distributions. Figure˜17 clearly shows that almost all the variance in the output at all locations is determined by $k_{ub}$.
@@ F.3 Sensitivity Analysis
F.4 Parameter Posterior for Remaining Parameters
Figure˜18 shows the KDEs of 4,000 parameter samples for the 3 remaining parameters, $c_{a,1},c_{a,2}$, and $\rho_{f}$, obtained with low-budget ABI, Point, E-Post, SABI, and UA-SABI. Two distinct scenarios emerge: 1) the posterior collapses to a point estimate (often at the boundary of the prior), or 2) the posterior remains close to the prior distribution. Point-like posteriors are produced by low-budget ABI or Point, while surrogate-based methods yield posteriors resembling the prior. Both outcomes can be explained by the lack of sensitivity of the output to these parameters, as shown by the total Sobol indices in Fig.˜17. This missing sensitivity makes reliable inference very difficult.
Figure 18: KDEs of 4,000 parameter posterior samples for the remaining parameters $c_{a,1},c_{a,2}$, and $\rho_{f}$ of the MICP model obtained with low-budget ABI, Point, E-Post, SABI, and UA-SABI.
F.5 Runtime Comparison